OpenStudy (annkisa):

Coulomb's constant-8,987,600,000N*m^2/C^2-is used to calculate the electric force between two stationary charged particles. Write the value for Coulomb's constant in scientific notation. Can someone please explain this to me?

OpenStudy (bugman954):

Write the value for Coulomb's constant in scientific notation \[-8.9876\times10^{9}\] or \[-8.99\times10^{9}\] if you are rounding

OpenStudy (vipulrakta):

SIMPLY THE COULOMBS CONSTANT IS \[1/4\pi \epsilon _{0}\] \[\epsilon _{0}= absolute \space permittivity \in free \space \] absolute permittivity is the measure of resistance that is encountered when forming an electric field in a medium. equal to 8.854 *10^-12 columbs constant = 1/[4*3.14*8.854*10*(-12)]=8.98* 10^9 N/m^2 c^2 \[k = 1/(4\pi \epsilon _{0}\epsilon) \space = 8.98* 10^{9} Nm ^{2}C ^{-2}\]

OpenStudy (annkisa):

Thank you both!